5-30 If there is a way to solve without excel and just using formulas please use that one.

74570445_547165619471008_283301408976299
Answer #1

5-30

Case of Foxhill instrument

Initial Cost=C=-$70000

Net Revenue per year=R=(Annual Income-AOC)=6500-5000=$1500

Salvage Value=S=$10000

NPW of option =C+R*(P/A,12%,15)+S*(F/P,12%,15)

NPW of option =-70000+1500*(P/A,12%,15)+10000*(P/F,12%,15)

Let us calculate interest factors

(P/F,0.12,15)=1/(1+0.12)^15=0.182696

1 (1+i) P/A, i, n)

1 (1+0.12)15 0.12 P/A, 0.12, 15) 6.810864

NPW of option =-70000+1500*6.810864+10000*0.182696=-57956.74

Case of Quicksliver

Initial Cost=C=-$50000

Net Revenue per year=R=(Annual Income-AOC)=5000-2500=$2500

Salvage Value=S=0

NPW of option =C+R*(P/A,12%,15)+S*(F/P,12%,15)

NPW of option =-50000+2500*(P/A,12%,15)+0*(P/F,12%,15)

NPW of option =-50000+2500*6.810864+0*0.182696=-32972.84

Case of Almaden

Initial Cost=C=-$90000

Net Revenue per year=R=(Annual Income-AOC)=7800-5000=$2800

Salvage Value=S=$10000

NPW of option =C+R*(P/A,12%,15)+S*(F/P,12%,15)

NPW of option =-90000+2800*(P/A,12%,15)+10000*(P/F,12%,15)

NPW of option =-90000+2800*6.810864+10000*0.182696=-69102.62

We can observe that NPW is highest (Least present cost) in case of Quicksilver. It should be purchased.